OpenStudy (anonymous):

Solve the initial-value problem y'+6xy=0;y(pi)=5

OpenStudy (anonymous):

\[\frac{dy}{dx}=-6xy \rightarrow dy*\frac{1}{y}=-6x dx\] \[\int\limits y^{-1}dy=-6 \int\limits x dx \implies \ln|y|=-3x^2+C\] \[\huge y=e^{-3x^2+c} \implies y=Ce^{-3x^2}\] \[\huge 5=Ce^{-3(\pi^2)} \implies 5e^{3\pi^2}=C\]

OpenStudy (anonymous):

I got that as well but the book says the answer is 5e^-3(x^2-pi^2)

OpenStudy (anonymous):

Oh, well then you do this: \[\huge y=Ce^{-3x^2}; C=5e^{3\pi^2} \implies y=5e^{3\pi^2}e^{-3x^2}\] \[\huge =5e^{-3(x^2-\pi^2)}\] Using the property that: \[(e^a)(e^b)=e^{a+b}\]